OpenStudy (anonymous):

Find the value of y so that so that the line passing through (4, y) and (5, 3) has a slope of 1.

ganeshie8 (ganeshie8):

\[slope = \frac{y2-y1}{x2-x1} = 1\] \[= > \frac{3-y}{5-4} = 1\] \[= > \frac{3-y}{1} = 1\] \[= > 3-y = 1\] \[= > y=2\]

OpenStudy (anonymous):

thank you @ganeshie8